### v1

# Do Robot Snakes Dream like Electric Sheep? Investigating the Effects of Architectural Inductive Biases on Hallucination

Abstract
The growth in prominence of large language models (LLMs) in everyday life can be largely attributed to their generative abilities, yet some of this is also owed to the risks and costs associated with their use. On one front is their tendency to hallucinate false or misleading information, limiting their reliability. On another is the increasing focus on the computational limitations associated with traditional self-attention based LLMs, which has brought about new alternatives, in particular recurrent models, meant to overcome them. Yet it remains uncommon to consider these two concerns simultaneously. Do changes in architecture exacerbate/alleviate existing concerns about hallucinations? Do they affect how and where they occur? Through an extensive evaluation, we study how these architecture-based inductive biases affect the propensity to hallucinate. While hallucination remains a general phenomenon not limited to specific architectures, the situations in which they occur and the ease with which specific types of hallucinations can be induced can significantly differ based on the model architecture. These findings highlight the need for better understanding both these problems in conjunction with each other, as well as consider how to design more universal techniques for handling hallucinations.
Do Robot Snakes Dream like Electric Sheep?
Investigating the Effects of Architectural Inductive Biases on Hallucination
[…]
1 Introduction
Large language models (LLMs) have rapidly emerged as a every-day tool in modern life (OpenAI, 2024), with many relying on their abilities to accomplish a variety of specific tasks. However, this opened up concerns relating to their propensity to hallucinate (Huang et al., 2023), with no concrete reasons for this behaviour (Dziri et al., 2022b; Rawte et al., 2023; Chen et al., 2024), hindering the ability to directly train LLMs that are consistently factual or able to explain themselves through their factual knowledge (Madsen et al., 2024a).
In parallel, as LLMs evolve and existing limitations are discovered, alternative architectures have become increasingly common and grow in popularity. In particular, Transformer LLMs (Vaswani et al., 2017) and linear sequence models (Gu et al., 2022; Gu and Dao, 2024) present contrasting methods of encoding sequences, with the Transformer using attention (Bahdanau et al., 2015) to form lossless representations of the context, while linear sequence models follow the recurrent neural network (Rumelhart et al., 1986; Jordan, 1986) in their use of a compressed state representation.
With the intensifying focus in both directions, a notable void exists in verifying how each can affect the others in conjunction. For example, existing works in hallucination detection and mitigation focus almost exclusively on Transformer-based models (Maynez et al., 2020; Longpre et al., 2021a; Guerreiro et al., 2023; Shi et al., 2023; Ji et al., 2023b; Farquhar et al., 2024; Wei et al., 2024), without extension to recurrent-style models, despite the use of a unified hidden representation potentially acting as an information bottleneck that can induce more common hallucinations. This lack of unified understanding on both topics prompts the need for a more explicit investigation.
In this work, we comprehensively explore the differences between pure attention LLMs and recurrent LLMs, specifically with respect to the propensity to hallucinate. Using a set of 20 different hallucination tasks, categorized into 6 groups that evaluate both faithfulness and factuality hallucinations, we evaluate across numerous open-source LLMs that range in scale from under 1B parameters to 70B parameters all the while covering a variety of different architecture choices such as self-attention, recurrent and hybrid models. We further evaluate across factors such as instruction-tuning, all to build a more comprehensive picture of architecture-specific phenomena with respect to hallucination. From this, we observe the following:
  1. (1)
    Viewed very broadly over various different tasks and settings, neither Transformer-based nor recurrent/hybrid LLMs appear to induce hallucinations more often than others.
  2. (2)
    However, shifting to individual tasks, it becomes evident that they result in disparate tendencies on specific tasks evaluating for unique criteria such as recalling long-tailed factual knowledge and falling into memorization traps, highlighting that model architecture may promote specific behavior that renders some types of hallucinations more common.
  3. (3)
    Evaluating the effects of scale and instruction-tuning and how these changes affect performance, we observe that though factuality hallucinations remain consistently dependent on model size, recurrent/hybrid architectures are often more faithful at smaller sizes and observe significantly fewer faithfulness benefits from instruction-tuning and scaling compared to self-attention models.
These results highlight that while some hallucinations may be quite consistent across architectures, others are the direct results of specific model design choices that go into the LLM construction. This hints towards the need for more careful consideration on this front, as different techniques for addressing this problem can potentially be highly catered towards specific models, bringing to the forefront the need for better consideration of both these problems in the face of each other.
2 Related Work
Hallucinations in LLMs.
Though various specific definitions exist, hallucination can broadly refer to "the generation of nonsensical information unfaithful to the provided source content" (Ji et al., 2023a). While this topic has become particularly important in order to ensure the safe and responsible use of language models, both classifying and quantifying the hallucinations is challenging. In particular, it is difficult to ascertain if the divergence occurs because of specific data heuristics (Lebret et al., 2016; Wiseman et al., 2017) or because of the innate lack of similarity between pre-training and downstream tasks (Rashkin et al., 2021), while measuring for hallucinations automatically introduces various biases (Reiter, 2018; Tian et al., 2020; Ganguli et al., 2022) that may not fully capture the scope of the errors. However, a variety of task-specific benchmarks (Li et al., 2020; Pagnoni et al., 2021; Zhou et al., 2021; Santhanam et al., 2022) have shown various LLMs to struggle with factual inconsistencies, highlighting a need to render them safer for every-day use.
LLM Architectures.
General large language models architecture can be broadly thought to be composed of two components: token mixers which serves to model transformations between time steps (such as attention and recurrent layers) and channel mixers, such as multi-layer perceptrons and mixtures-of-experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017), which allow communication between different channels within a single time step. Accordingly, token mixing often forms a computational bottleneck in terms of time complexity while channel mixers consist of a memory bottleneck. Though the contemporary standard for token mixing remains self-attention, alternatives are becoming increasingly common as they begin to display more promise. These include the use of linear attention (Katharopoulos et al., 2020) to compensate for the quadratic memory complexity of vanilla self-attention, to new recurrent models (Gu and Dao, 2024; Team, 2024b) that function as a linear recurrent neural network but can process all elements of a sequence in parallel, and well as hybrid mixtures of recurrent mechanisms and attention (AI21, 2024; Dao and Gu, 2024) that have emerged as a meaningful competitor.
Architecture and Hallucination.
Despite the growing research in new architectural components, whether or not how they affect hallucination has yet to be studied. While some works (Madsen et al., 2024b; Hu et al., 2024; Schimanski et al., 2024) have proposed modifications to either the learning/generation pipeline as a way of reducing hallucinations, explicit proposals for hallucination reduction through structural modification have yet to be suggested. Additionally, though some works (Elhage et al., 2021; Fu et al., 2023; Lutati et al., 2023; Poli et al., 2024) have demonstrated self-attention to aptly solve synthetic tasks that form an essential component of language modelling, these fail to remain faithful on more realistic datasets, punctuating a major limitation of existing LLMs. Finally, though tangential work demonstrates that recurrent models may suffer from issues with information retention (Vardasbi et al., 2023), formally defining a link with hallucination remains necessary. These issues lead to a lack of clarity on this front, accentuating the need for a formal investigation comparing hallucination alongside architectural paradigms. In particular, the choice of token mixer is important, as this is the component that governs how information is shared between different elements of the sequence to form a holistic representation. As such, an effective token mixer will adequately enable enough information from the context to propagate forward during the generation phase (Olsson et al., 2022; Arora et al., 2024), potentially enabling models to hallucinate less.
3 Background
Attention for Sequences.
Vanilla self-attention as used in Transformers is powerful but costly. When provided an embedded text representation as a sequence of tokens $\bm{x}\in\mathbb{R}^{L\times d}$, each Transformer layer in the network applies a function
$T_{\ell}(\bm{x})=\text{FF}_{\ell}(A_{\ell}(\bm{x})+\bm{x})+A_{\ell}(\bm{x})$ (1)
where $A_{\ell}$ is the self-attention of the $\ell$-th layer and $\text{FF}_{\ell}$ is the following feed-forward network. Self-attention computes, for a token at position $i$ in a sequence, a weighted average of the feature representations of all tokens (the values $\bm{V}_{\ell}$) in the sequence with a weight proportional to a similarity score between $i$ (the query $\bm{Q}_{\ell}$ at position $i$) and the rest of the sequence (the keys $\bm{K}_{\ell}$). In particular, these can be computed for all positions in parallel
$\begin{split}&\bm{Q}_{\ell}=\bm{x}\bm{W}_{\ell}^{\bm{Q}}\quad\bm{K}_{\ell}=\bm{x}\bm{W}_{\ell}^{\bm{K}}\quad\bm{V}_{\ell}=\bm{x}\bm{W}_{\ell}^{\bm{V}}\\ &A_{\ell}(\bm{x})=\bm{V}^{\prime}_{\ell}=\text{softmax}\big{(}{\bm{Q}_{\ell}\bm{K}_{\ell}^{T}}/{\sqrt{d}}\big{)}\bm{V}_{\ell}\end{split}$ (2)
providing the model a lossless representation of the complete past context. This can be seen as equivalent to search and retrieval within a database, where search is defined using query-key parameterizations and retrieval with value parameterization. In this setting, the database from which the information is being retrieved is equivalent to the model parameters, which store information from a training corpus that the weights $\bm{W}^{\bm{K}},\bm{W}^{\bm{Q}},\bm{W}^{\bm{V}}$ parameterizing the model attempt to mimic.
Multi-head attention, a variant of self-attention popularized by the Transformer (Vaswani et al., 2017), has become the dominant variant used in LLMs such as LLaMA (Touvron et al., 2023) and Gemma (Team, 2024a). Additional variants, such as multi-query attention from Falcon (Almazrouei et al., 2023) and grouped-query attention from Mistral (Jiang et al., 2023), adapt multi-head attention but remain build upon the same underlying principle of self-attention. Consequently, such methods are considered to fall under the family of self-attention token mixers.
From Transformers to Recurrent LLMs.
One of the original concerns of Transformers was the quadratic complexity of attention, leading to large amounts of work focused on improving this bound as it could directly affect the ability to learn from long sequences. Instead of directly modifying attention (Katharopoulos et al., 2020; Kitaev et al., 2020; Choromanski et al., 2021), Gu et al. (2020) motivated a novel paradigm using state-space models (SSMs) from control theory. SSMs map an input $x(t)\in\mathbb{R}^{d}$ to an intermediate state $h(t)\in\mathbb{R}^{n}$ that is then projected to an output $y(t)\in\mathbb{R}^{d}$:
$h^{\prime}(t)={\bm{A}}h(t)+{\bm{B}}x(t),\quad y(t)={\bm{C}}h(t)+{\bm{D}}x(t)$
where $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ are all trainable parameters. Gu et al. (2021) use this paradigm to define a recurrent model to work on discrete signals, in which case the input can be regarded as discretized data sampled from a continuous signal with a step size $\Delta$, for which the corresponding SSM is defined by:
$\begin{split}h_{t}&=\overline{\bm{A}}h_{t-1}+\overline{\bm{B}}x_{t}\quad y_{t}=\overline{\bm{C}}h_{t}+\overline{\bm{D}}x_{t}\\ \overline{\bm{A}}&=\frac{\big{(}I+{\Delta}\bm{A}/{2})}{\big{(}I-{\Delta}\bm{A}/{2}\big{)}}\quad\overline{\bm{B}}=\frac{\Delta\bm{B}}{\big{(}I-{\Delta}\bm{A}/{2}\big{)}}\end{split}$
and $\overline{\bm{C}}=\bm{C}$ ($\overline{\bm{D}}$ is equivalent to a residual connection and set to $\bm{0}$.) Thus
$\begin{split}\overline{\bm{K}}&=(\overline{\bm{CB}},\overline{\bm{CAB}},\dots,\overline{\bm{CA}}^{L-1}\overline{\bm{B}})\quad\bm{y}=\overline{\bm{K}}*\bm{x}\end{split}$
where $\overline{\bm{K}}$ is the SSM kernel. As $\bm{y}$ can be computed in $O(L\log L)$ with a Fast Fourier Transform (Cormen et al., 2009), the entire output can be computed in tandem based on the input, given the matrices that parameterize the system. Furthermore, setting $\bm{A}$ as a Hurwitz matrix, SSMs can preserve long-term dependency information and enable them to capture long-dependence information, overcoming a long-standing issue (Bengio et al., 1994) with prior recurrent models (Hochreiter and Schmidhuber, 1997; Cho et al., 2014).
Future works then modify this general structure for additional properties; Mamba (Gu and Dao, 2024) use input-dependent $\bm{B}$ and $\bm{C}$, enabling for the model to filter information. LRU/Hawk (Orvieto et al., 2023; De et al., 2024) remove the discretization step from a continuous signal and instead learn discrete matrices for $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ directly. RWKV models (Peng et al., 2023, 2024) use a novel WKV token mixing operator that acts like an RNN. However, the token mixing paradigm for all such models shares one commonality, namely the use of hidden states from which token-level information must pass through in order to reach another token further in the future. Further interest in this topic has led to the creation of hybrid mixtures of recurrence and attention (Dao and Gu, 2024; De et al., 2024; AI21, 2024) to form new classes of models meant to capture the benefits of both.
Different time-mixing modules attempt to encode the history of tokens in a fixed representation. Recurrent architectures maintain this fixed representation and learn to update it during pre-training. Despite the variety of mechanisms and improvements to the representation according to the design of the module, they can fall short of attention (Vardasbi et al., 2023; Jelassi et al., 2024). One reason is that the amount of information that can be encoded (irrespective of length) is constrained by the size of hidden state (Arora et al., 2024). However, available architectures all converge during training on their respective pre-training and achieve comparable metrics. This then suggests that each architecture learns different aspects of the sequences that align with its own inductive biases. Self-attention architectures have the luxury to look at long range dependencies (albeit being costly). On the other hand, recurrent structures are constrained to make the predictions based on the recent context (which in practice can be long, but remains shorter than what self-attention can attend to). But the expressivity of the hidden representation as modeled by the time-mixers directly affects its performance on tasks (Sun et al., 2024) and the prior knowledge in the parameters of the architectures can be predisposed to favor the prediction of certain tokens. We model this effect on tasks under the lens of hallucination: faithfulness and factuality. Faithfulness expects the model to follow an instruction (that is largely recent), while factuality assumes the use of token-to-token relations learned during pre-training for a broader contextual representation. Following this, we ask two questions: (1) Do self-attention and recurrent architectures hallucinate differently? (2) Does instruction-tuning and/or size benefit the different classes of LLMs equally?
Figure 1: Performance on PopQA (left) and MemoTrap (right). Recurrent and hybrid models significantly outperform similar pure attention-based alternatives on MemoTrap, but the opposite is true on PopQA.
4 Experiments and Results
Models.
For our comparison, we use various models (Table 2) that vary in size and architecture, with a particular focus on different time-mixing methods. These include self-attention models such as Pythia, LLaMA2/3 (Touvron et al., 2023), Falcon-7B (Almazrouei et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma1/2 (Team, 2024a) and Mixtral-8x7B (Jiang et al., 2024). We also test recurrent models including Mamba (Gu and Dao, 2024; Dao and Gu, 2024), FalconMamba-7B and RWKV (Finch) (Peng et al., 2023), as well as hybrid models like RecurrentGemma (Team, 2024b) and Jamba (AI21, 2024). For models with instruction-tuned versions, we use both the base and instruction-tuned variant. Additional details are available in Appendix A.
Datasets.
We evaluate on the Hallucination Leaderboard (Hong et al., 2024), consisting of tasks
  1. 1)
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (MC1, MC2, Generative) (Lin et al., 2022), PopQA (Joshi et al., 2017)
  2. 2)
    Summarization: XSum (Narayan et al., 2018), CNN/DM (See et al., 2017)
  3. 3)
    Reading Comprehension: RACE (Lai et al., 2017), SQuADv2 (Rajpurkar et al., 2018), NQ-Swap (Longpre et al., 2021b)
  4. 4)
    Instruction Following: MemoTrap (Liu and Liu, 2023), IFEval (Zhou et al., 2023)
  5. 5)
    Hallucination Detection: FaithDial (Dziri et al., 2022a), HaluEval (QA, Summarization, Dialogue) (Li et al., 2023a), True-False (Azaria and Mitchell, 2023)
  6. 6)
    Fact Checking: FEVER (Thorne et al., 2018)
[…]
4.1 Investigating Task Biases
At a broader scope that architecture does not appear to influence hallucination, both for task categories and factuality/faithfulness. For example, with like-sized models falling within different categories of their time mixers (self-attention, recurrent layers, hybrid), mean performance on tasks within each grouping is not significantly different nor do they show a particular ordering [^1] Explicit numbers in Table 3, 4 and 5 in Appendix B.. However, with specific tasks, differences start to present themselves, which we describe in our following observations.
O1: Recurrent models struggle with rare knowledge.
Interestingly, recurrent and hybrid models all significantly underwhelm on PopQA, a task that tests for uncommon (thin-tailed) factual knowledge (left plot in Figure 1). For example, FalconMamba achieves an exact-match (EM) score of 0.7 compared to 17.5 by a similar Falcon. Similarly, 2B and 9B RecurrentGemma achieve 7.6 and 16.0 EM respectively, compared to 14.6/21.3 EM for 2B/7B Gemma and 13.6/18.2 by 2B/9B Gemma2, while Jamba (0.4 EM) performs worse than a similarly sized Mixtral-8x7B (31.9 EM).
Previous work (Vardasbi et al., 2023) suggests that hidden states can become dense and difficult to extract information from in the future. Although this is relevant for hallucinations, these concerns are unlikely to apply in this setting given that the prompt consists of only a question and the amount of information to store within the hidden representation is unlikely to lead to it becoming highly dense. Additional works (Dao and Gu, 2024) have instead posited that using recurrent models with additional mechanisms enabling input filtering can enhance performance on information-dense data such as language; however, this is again unlikely to be the case due to the general simplicity of the task. Yet this result is significant as it indicates recurrent models, despite the task’s simplicity, struggle when compared to attention-based equivalents, suggesting a direct link to the token mixing that is utilized. The reason may instead be that recurrent/hybrid models do not learn by explicitly memorizing and acting as a retrieval system (as discussed in Section 3). In this sense, despite potentially observing low-tail information, the learning mechanism may not have led to said information being stored within parameters, making recurrent/hybrid models more likely to produce incorrect information in such settings with higher degrees of uncertainty.
Model Name Hallu. Detection Instr. Following Closed-Book QA Reading Comp. Sum. Fact-Checking Faithfulness Factuality
Attention-Only Models
Gemma-2B 49.65 ($\uparrow$ 11.51) 30.25 ($\uparrow$ 13.84) 38.20 ($\downarrow$ -6.92) 38.96 ($\downarrow$ -13.96) 24.93 ($\downarrow$ -4.53) 54.14 ($\uparrow$ 1.60) 35.85 ($\uparrow$ 1.92) 32.56 ($\downarrow$ -4.56)
Gemma-7B 53.81 ($\downarrow$ -5.67) 31.94 ($\uparrow$ 6.79) 27.43 ($\downarrow$ -4.13) 31.35 ($\downarrow$ -5.83) 20.63 ($\uparrow$ 1.04) 43.28 ($\downarrow$ -5.22) 36.48 ($\uparrow$ 2.71) 44.50 ($\downarrow$ -6.04)
Gemma2-2B 58.33 ($\uparrow$ 0.64) 26.73 ($\uparrow$ 10.90) 29.59 ($\uparrow$ 4.83) 32.95 ($\downarrow$ -5.26) 16.04 ($\downarrow$ -0.47) 39.57 ($\uparrow$ 18.96) 35.61 ($\uparrow$ 0.65) 34.96 ($\downarrow$ -0.98)
Gemma2-9B 65.64 ($\uparrow$ 6.31) 23.76 ($\uparrow$ 30.07) 39.36 ($\uparrow$ 3.13) 39.74 ($\downarrow$ -0.70) 21.38 ($\downarrow$ -3.35) 62.06 ($\uparrow$ 6.47) 40.56 ($\uparrow$ 7.55) 46.85 ($\uparrow$ 3.13)
Gemma2-27B 62.03 ($\uparrow$ 12.43) 26.62 ($\uparrow$ 33.58) 47.19 ($\uparrow$ 2.86) 43.04 ($\uparrow$ 5.02) 28.92 ($\downarrow$ -0.74) 68.28 ($\uparrow$ 1.00) 41.54 ($\uparrow$ 13.02) 53.92 ($\uparrow$ 2.36)
LLaMA2-7B 53.63 ($\uparrow$ 4.30) 28.94 ($\uparrow$ 8.99) 37.64 ($\downarrow$ -1.89) 27.58 ($\uparrow$ 3.74) 25.19 ($\uparrow$ 0.69) 51.38 ($\uparrow$ 5.88) 35.14 ($\uparrow$ 4.01) 42.51 ($\downarrow$ -0.12)
LLaMA2-13B 67.80 ($\downarrow$ -3.68) 26.57 ($\uparrow$ 9.72) 39.65 ($\downarrow$ -0.65) 32.17 ($\downarrow$ -3.64) 27.33 ($\downarrow$ -0.23) 62.35 ($\downarrow$ -1.64) 40.83 ($\uparrow$ 0.88) 46.27 ($\downarrow$ -0.49)
LLaMA2-70B 62.34 ($\uparrow$ 10.78) 30.24 ($\uparrow$ 17.11) 47.85 ($\downarrow$ -2.53) 39.48 ($\downarrow$ -8.21) 28.00 ($\downarrow$ -0.55) 66.63 ($\downarrow$ -1.45) 41.74 ($\uparrow$ 5.82) 53.69 ($\downarrow$ -2.34)
LLaMA3-8B 60.76 ($\uparrow$ 10.12) 22.06 ($\uparrow$ 26.76) 41.55 ($\uparrow$ 1.30) 33.52 ($\uparrow$ 3.62) 26.62 ($\downarrow$ -1.60) 60.84 ($\uparrow$ 4.82) 37.60 ($\uparrow$ 10.09) 48.14 ($\uparrow$ 1.60)
LLaMA3-70B 71.78 ($\uparrow$ 8.69) 21.59 ($\uparrow$ 25.60) 48.07 ($\uparrow$ 3.43) 46.38 ($\downarrow$ -7.73) 28.91 ($\downarrow$ -1.68) 69.57 ($\uparrow$ 0.99) 45.92 ($\uparrow$ 6.23) 54.80 ($\uparrow$ 2.81)
Mistral-7B 60.48 ($\uparrow$ 3.36) 28.39 ($\uparrow$ 16.85) 41.24 ($\uparrow$ 5.21) 32.03 ($\uparrow$ 1.92) 26.77 ($\downarrow$ -0.77) 58.59 ($\uparrow$ 6.78) 38.47 ($\uparrow$ 4.67) 47.42 ($\uparrow$ 4.17)
Mixtral-8x7B 73.51 ($\downarrow$ -0.15) 26.84 ($\uparrow$ 17.60) 48.99 ($\uparrow$ 4.33) 40.66 ($\downarrow$ -3.90) 27.91 ($\downarrow$ -1.07) 68.35 ($\uparrow$ 0.49) 46.16 ($\uparrow$ 3.89) 55.15 ($\uparrow$ 3.50)
Falcon-7B 52.40 ($\downarrow$ -2.90) 25.54 ($\uparrow$ 12.62) 33.03 ($\downarrow$ -4.42) 29.18 ($\downarrow$ -3.64) 20.70 ($\downarrow$ -1.00) 46.91 ($\downarrow$ -7.71) 34.91 ($\downarrow$ -0.71) 36.92 ($\downarrow$ -3.75)
Recurrent and Hybrid Models
RecurrentGemma-2B 52.88 ($\uparrow$ 2.33) 33.43 ($\uparrow$ 3.12) 27.36 ($\downarrow$ -1.85) 30.66 ($\downarrow$ -1.77) 18.41 ($\uparrow$ 2.86) 42.97 ($\uparrow$ 3.63) 36.55 ($\downarrow$ -0.14) 31.46 ($\uparrow$ 1.39)
RecurrentGemma-9B 55.75 ($\downarrow$ -1.67) 31.67 ($\uparrow$ 12.96) 36.79 ($\uparrow$ 2.03) 36.57 ($\downarrow$ -6.66) 22.99 ($\uparrow$ 1.84) 51.25 ($\uparrow$ 13.36) 37.61 ($\downarrow$ -0.17) 42.96 ($\uparrow$ 3.28)
Jamba 57.66 ($\downarrow$ -2.90) 43.36 ($\downarrow$ -5.76) 39.50 ($\downarrow$ -0.98) 33.00 ($\downarrow$ -5.88) 23.72 ($\downarrow$ -15.48) 59.88 ($\downarrow$ -2.60) 39.80 ($\downarrow$ -6.91) 45.95 ($\downarrow$ -0.91)
FalconMamba-7B 55.80 ($\uparrow$ 0.73) 42.97 ($\downarrow$ -1.19) 39.94 ($\downarrow$ -0.43) 23.76 ($\downarrow$ -0.85) 23.89 ($\downarrow$ -0.09) 61.85 ($\downarrow$ -1.80) 35.92 ($\downarrow$ -0.26) 47.02 ($\downarrow$ -0.38)
Table 1: Changes in performance from the use of instruction-tuning. While factuality does not exhibit a specific trend, faithfulness is shown to improve within attention-based models but not recurrent or hybrid models.
O2: Recurrent models can better follow contexts by relying less on memory.
There are also cases where recurrent models consistently outperform attention equivalents, namely on MemoTrap (right plot in Figure 1). Here, the LLM is prompted to complete a well-known proverb with an ending that deviates from the commonly used ending, testing for over-reliance on knowledge memorized from a training corpus. This suggests that although recurrent models might struggle with long-tail knowledge, they are less prone to ignoring contextual cues not stored within parametric memory, or they more model the dynamics based on the context as opposed to acting as a retrieval system. Noticeable drops exist from FalconMamba (58.8) to Falcon (38.9) and from RecurrentGemma 2B/9B to Gemma2 2B/9B (53.7/43.4 vs 42.7/34.4), suggesting that attention can lead to more hallucinations. It further reveals a benefit of recurrent layers; by not memorizing information directly within parameters, there may be a reduced tendency to repeat previous observations (Jelassi et al., 2024) and greater focus on the context.
O3: Scale is required for emergent qualities.
However, some scale to the data and model remain important. Note that Mamba, Finch and Pythia fail to draw the same distinctions as the other models despite sharing the same vocabulary and training data. However, given the scale of training for these (300B tokens compared to >1T tokens for other models) and the fact that the discussed tasks relate directly to information memorization, it is possible that these phenomena fail to emerge at this data scale. This is particularly evidenced by how these models perform better with size on MemoTrap, which is designed in a way such that larger models should normally perform worse.
4.2 On the role of Instruction-Tuning.
The use of instruction-tuning appears to show inconsistent improvements across categories (Table 1) and tasks [^2] See Figure 5.. There exists no strict pattern for overall hallucination; on some tasks, its use is effective for all (ex. TruthfulQA) or no (ex. TriviaQA) models. Smaller trends also exist, such as instruction-following observing the strongest gains.
However, changes in faithfulness are generally positive for pure attention models (shown in teal) and negative for models with recurrent layers (shown in magenta), whereas changes in factuality show no consistent trends. This suggests that instruction-tuning for LLMs, as initially proposed with self-attention based models, enables particular properties which are no longer relevant after the addition of recurrent layers. Their inductive biases therefore can have a particular effect on specific forms of hallucinations emerging, with instruction-tuning being a manner to mitigate these.
Recalling the links between attention and retrieval from Section 3, the process of instruction-tuning can be understood as shifting the projection weights towards a new space through the additional fine-tuning (which is of itself simply an additional step of language modeling). This leads to the over-writing of some of the pre-training corpus with that used for fine-tuning, effectively replacing the keys and values from the underlying retrieval storage. While effects are ambiguous on factuality, as some potentially relevant facts can be removed, faithfulness often improves from this process, as the fine-tuning corpus is more likely to contain information relevant to context following and thereby make it easier for the attention mechanism to retrieve relevant information for such a purpose. In addition to highlighting how enhancing the ability to follow instructions does not necessarily lead to factually correct results, another conclusion stemming from the fact that recurrent models generally observe no benefits in faithfulness after instruction-tuning is that such models’ inductive biases render them inherently more faithful or the process of making them more faithful is distinct from self-attention.
Figure 2: Changes in factuality (top) and faithfulness (bottom) as models are increased in size. Score range between 0 and 100. Factuality always increases with the number of parameters, however faithfulness increases are only meaningful for pure-attention models.
Figure 3: Performance to scale values for factuality and faithfulness. Colors differentiate model families while shapes differentiate the model type. Factuality improves with model size (indicated by values >0) and the model type does not have a link with this relative improvement. Recurrent/hybrid models show low values for faithfulness, indicating that size increases generally do not benefit them, unlike attention models.
4.3 Impact of Model Size on Hallucinations.
With the large amount of prior work (Kaplan et al., 2020; Wei et al., 2022; Hoffmann et al., 2022) concluding that model size can play a significant role in model reasoning abilities, we provide this additional axis of variation in Figure 2.
Immediately, we observe that increased size leads to improvements in factuality regardless of model. This is partially expected, considering the range of work that has suggested links between size and the amount of factual knowledge that can be encoded within parametric space. However, faithfulness is different. Recurrent and hybrid architectures tend to saturate in terms of faithfulness at a small size whereas pure attention models do not. For example, Mamba models stagnate in faithfulness at around 130M, whereas Pythia continues to increase in this score up to billions of parameters. Finch and RecurrentGemma also show this while attention-based models do not. Another observation is the fact that even the smallest Mamba are nearly as faithful as LLaMA2-7B and Pythia-1.4B and despite models eventually catching up in terms of faithfulness, having such small models perform on par indicates a meaningful manner in which their hallucinations may differ. To further illustrate this, we provide an additional depiction of a performance to scale metric, defined as
[…]
where $m(\cdot)$ is a metric of interest, $\mathcal{M}$ is a model and $\mathcal{B}$ is a base model of the family as $\mathcal{M}$. $P(\cdot)$ counts the number of parameters in a given model. Namely, this measures the relative increase in $m$ by linear scaling the model, with a lower value indicating that linearly increasing the number of parameters yields no changes while larger values indicate scaling leads to significant improvements. More positive values indicate that the score increases greatly with linear increases in size and negative values the opposite. Consistently negative or near-zero values hence mean that size has little effect on the score. Figure 3 depicts the performance to scale changes for factuality and faithfulness, where it becomes apparent that increasing the number of parameters in hybrid and recurrent models yields a significantly lower increase in faithfulness compared to pure attention models.
This suggests that recurrent layers enable an inherent ability to follow contexts not improved by model size. Their inductive biases can be perceived as encouraging context following, supported by how the hidden state incorporates information from the context with the current input to control the generation. This perspective aligns these results with our instruction-tuning observations, where such models do not exhibit a clear benefit compared to their attention-based counterparts. In particular, we can interpret this as indicating that the base model learns to perform this output control even without direct instruction-tuning; hence applying these techniques is ineffective as we previously observed.
5 Discussion
Hallucination Mitigation for Different Architectures.
Hallucinations motivate the need to both detect and mitigate them, particularly when they can escalate to have adverse effects. Existing techniques may be insufficient, as they contain potential implicit biases that can fundamentally conflict with the manners in which hallucinations may differ between models. Facts such as sequence models saturating in faithfulness in a manner not resolved through instruction-tuning indicate that existing mitigation techniques can differ in their effectiveness due to a bias towards specific architectures.
This means that current techniques also further deserve a look, in particular in terms of their effectiveness across the different inductive biases of models. While some methods evidently might be specifically catered to specific models, e.g. using attention (Li et al., 2023b; Zhang et al., 2024; Chuang et al., 2024), other methods may implicitly also possess biases due to the underlying assumptions being made or specific requirements that may be unevenly addressed through different architectures, such as the reliance on data-refinement or additional context (Shi et al., 2024).
Inherent Faithfulness and Factuality.
Designing inherently faithful (Herman, 2019; Wiegreffe and Pinter, 2019) or factual models remains an important issue. While much work suggests attention-based LLMs lack an inherent tendency to be faithful (Jacovi and Goldberg, 2020; Wiegreffe and Marasovic, 2021), methods have been designed to render models more interpretable (Madsen et al., 2024b) for specifically evaluating this. Similarly, some of these methods have been shown to also adapt to Mamba models (Sharma et al., 2024). However, while interpretability is useful for evaluating models as being faithful or factual, it does not ensure an inherent tendency towards being either.
Nevertheless, these results hint that architectures can provide an avenue through which desiderata can be attained. SSMs present an option which in particular appears suited for settings which require utilizing information provided within a context, given their use within control theory to model system dynamics. With language, this can be analogous to using the fixed-size representation of the previous context for control, which may mean that such models can remain more faithful as the context remains the source that guides the generation. In comparison, attention-based models have been shown to act as retrieval systems that retrieve tokens based on high similarity with memorized information; hence, when a large difference exists between the context and what has been stored from training, breakdowns in faithfulness can occur.
6 Conclusion
Are hallucinations the direct result of how token-mixing takes place across time? We study whether different inductive biases of LLM architectures can increase this propensity by observing how hallucinations change across numerous tasks. Patterns emerge demonstrating that some architectural biases can lead to either improvements or degradation in performance compared to others, both for individual tasks and categorizations. Additionally, some model types lead to inherent behaviour when comparing the effects of instruction-tuning and scaling model size. In sum, model-specific inductive biases can have a direct effect on the type of hallucinations they are faced with. However, the data used and the specific model construction can also affect learning. We hope that future work can build on our findings by exploring both the types of hallucinations that can occur and techniques meant to mitigate them and how to design more universal techniques to make models more robust or reliable.
7 Limitations
Evaluating faithfulness and factuality.
A major limitation existing in hallucination detection and mitigation research is the lack of metrics that provide explicit information regarding hallucination. While the development of such metrics is an active area of research, limitations still exist, such as biases when involving additional models within the evaluation process.
Limits on tasks.
Another limitation of this work is the non-exhaustive set of tasks and domains in which we measure hallucinations. Due to the exhaustive ways in which hallucinations can be measured, we limit ourselves to tasks that are well motivated and frequently used in practice.
8 Ethical Concerns
This paper provides an analysis on the effects of different architectural inductive biases on the propensity to hallucinate within large language models. As such, mistakes in methodology can lead to unsupported confidence or skepticism regarding their performance with the explored task or related ones. While skepticism may not be an ethical issue, unsupported confidence can be problematic. However, the overall message is that all LLMs have specific tasks and settings in which they will display a greater propensity to hallucinate, hence we do not believe that the ideas expressed in this work will explicitly lead to unsupported confidence.
9 Acknowledgements
Jerry Huang received partial support under the forms of a National Science and Engineering Research Council (NSERC) Canada Graduate Scholarship, a Fonds de Recherche du Québec Nature et Technologies (FRQNT) Training Scholarship, and a Hydro-Québec Excellence Scholarship over the course of this work. Sarath Chandar is supported by a Canada CIFAR AI Chair, the Canada Research Chair in Lifelong Machine Learning, and a NSERC Discovery Grant.
Appendix A Technical Implementation Details
A.1 Models and Datasets
All models and datasets used in this paper are public and directly available on the HuggingFace Hub.
A.2 Experimental Setup
Experiments were run using the Language Model Evaluation Harness (Gao et al., 2024).
A.3 Datasets
We evaluate on the Hallucination Leaderboard (Hong et al., 2024), consisting of tasks
  1. 1)
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (MC1, MC2, Generative) (Lin et al., 2022), PopQA (Joshi et al., 2017)
  2. 2)
    Summarization: XSum (Narayan et al., 2018), CNN/DM (See et al., 2017)
  3. 3)
    Reading Comprehension: RACE (Lai et al., 2017), SQuADv2 (Rajpurkar et al., 2018), NQ-Swap (Longpre et al., 2021b)
  4. 4)
    Instruction Following: MemoTrap (Liu and Liu, 2023), IFEval (Zhou et al., 2023)
  5. 5)
    Hallucination Detection: FaithDial (Dziri et al., 2022a), HaluEval (QA, Summarization, Dialogue) (Li et al., 2023a), True-False (Azaria and Mitchell, 2023)
  6. 6)
    Fact Checking: FEVER (Thorne et al., 2018)
[…]
A.4 Models and Baselines
For our comparison, we use various models (Table 2) that vary in size and architecture, with a particular focus on different time-mixing methods. These include:
  • All Pythia models
  • All publicly available Mamba (Gu and Dao, 2024; Dao and Gu, 2024) models
  • All LLaMA2 and LLaMA3 (Touvron et al., 2023) models
  • Falcon-7B (Almazrouei et al., 2023) and FalconMamba-7B
  • Mistral-7B (Jiang et al., 2023)
  • All Gemma (Team, 2024a) models
  • RWKV (Peng et al., 2023) models, namely Finch (RWKV-v6)
  • RecurrentGemma (Team, 2024b)
  • Mixtral-8x7B (Jiang et al., 2024)
  • Jamba (AI21, 2024)
For models with instruction fine-tuned versions, we use both the base version as well as the instruction fine-tuned variant.
Model Public Link HuggingFace Model
Pythia-160M EleutherAI/pythia-160m
Pythia-410M EleutherAI/pythia-410m
Pythia-1B EleutherAI/pythia-1b
Pythia-1.4B EleutherAI/pythia-1.4b
Pythia-2.8B EleutherAI/pythia-2.8b
Pythia-6.9B EleutherAI/pythia-6.9b
Pythia-12B EleutherAI/pythia-12b
Mamba-130M state-spaces/mamba-130m
Mamba-370M state-spaces/mamba-370m
Mamba-790M state-spaces/mamba-790m
Mamba-1.4B state-spaces/mamba-1.4b
Mamba-2.8B state-spaces/mamba-2.8b
Mamba2-130M state-spaces/mamba2-130m
Mamba2-370M state-spaces/mamba2-370m
Mamba2-780M state-spaces/mamba2-780m
Mamba2-1.3B state-spaces/mamba2-1.3b
Mamba2-2.7B state-spaces/mamba2-2.7b
Mamba2Attention-2.7B state-spaces/mamba2attn-2.7b
RecurrentGemma-2B google/recurrentgemma-2b
RecurrentGemma-2B (IT) google/recurrentgemma-2b-it
RecurrentGemma-9B google/recurrentgemma-9b
RecurrentGemma-9B (IT) google/recurrentgemma-9b-it
Gemma-2B google/gemma-2b
Gemma-2B (IT) google/gemma-2b-it
Gemma-9B google/gemma-9b
Gemma-9B (IT) google/gemma-9b-it
Gemma2-2B google/gemma2-2b
Gemma2-2B (IT) google/gemma2-2b-it
Gemma2-9B google/gemma2-9b
Gemma2-9B (IT) google/gemma2-9b-it
Gemma2-27B google/gemma2-27b
Gemma2-27B (IT) google/gemma2-27b-it
Falcon-7B tiiuae/falcon-7b
Falcon-7B (IT) tiiuae/falcon-7b-instruct
FalconMamba-7B tiiuae/falcon-mamba-7b
FalconMamba-7B (IT) tiiuae/falcon-mamba-7b-instruct
Mistral-7B mistralai/Mistral-7B-v0.3
Mistral-7B (IT) mistralai/Mistral-7B-Instruct-v0.3
Mixtral-8x7B mistralai/Mixtral-8x7B-v0.1
Mixtral-8x7B (IT) mistralai/Mixtral-8x7B-Instruct-v0.1
Jamba ai21labs/Jamba-v0.1
Jamba (IT) ai21labs/AI21-Jamba-1.5-Mini
LLaMA2-7B meta-llama/Llama-2-7b-hf
LLaMA2-7B (IT) meta-llama/Llama-2-7b-hf
LLaMA2-13B meta-llama/Llama-2-13b-hf
LLaMA2-13B (IT) meta-llama/Llama-2-13b-hf
LLaMA2-70B meta-llama/Llama-2-70b-hf
LLaMA2-70B (IT) meta-llama/Llama-2-70b-hf
LLaMA3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B
LLaMA3-8B (IT) https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
LLaMA3-70B https://huggingface.co/meta-llama/Meta-Llama-3-70B
LLaMA3-70B (IT) https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct
Table 2: Models used and public links to their weights.
A.5 Computing Resources Used
All results were obtained using a server of 8 NVIDIA V100 32GB GPUs. The accelerate package was used for model sharding in instances where a single GPU was insufficient to store the entire model.
Appendix B Complete Results
B.1 Numerical Results
Numerical results are provided in the following tables
  • Figure 4 and Table 3 contains results for all individual tasks.
  • Table 4 contains results on the different task categories.
  • Table 5 contains results for factuality and faithfulness.
Figure 4: Performance of various base models on tasks within the Hallucination Leaderboard.
Model HaluEval Dial. HaluEval QA HaluEval Sum. Memo-Trap PopQA TruthfulQA-Gen TruthfulQA-MC1 TruthfulQA-MC2 RACE IfEval CNNDM XSum SQuADv2 NQ (8-shot) TriviaQA (8-shot) FaithDial TrueFalse Fever10 NQ (64-shot) TriviaQA (64-shot) Avg
EM EM EM Accuracy EM Rouge-L Accuracy Accuracy Accuracy Accuracy Rouge-L Rouge-L EM Rouge-L Rouge-L Rouge-L Rouge-L Rouge-L EM EM
Falcon-7B 41.81 46.57 48.95 38.89 17.46 24.97 22.40 34.26 37.42 12.20 20.59 20.81 26.01 22.88 58.68 66.60 58.08 46.91 24.32 59.27 35.87
Falcon-7B (IT) 38.40 29.64 48.65 60.79 8.68 38.07 29.01 44.12 37.22 15.53 20.43 18.97 20.68 14.29 39.21 67.14 63.66 39.20 15.51 39.95 33.71
FalconMamba-7B 49.93 39.54 42.53 58.76 0.74 49.20 37.82 53.40 42.97 27.17 25.09 22.68 24.36 25.96 64.13 58.18 88.84 61.85 27.84 60.44 41.21
FalconMamba-7B (IT) 50.14 40.16 39.41 56.20 0.18 49.08 37.70 53.23 42.58 27.36 24.79 22.80 24.53 26.34 63.45 62.67 90.29 60.05 27.65 58.45 40.89
Finch-1.6B 47.72 51.05 50.32 45.94 1.58 26.07 22.15 36.73 35.60 16.27 6.63 16.91 10.97 10.30 36.71 63.24 53.10 39.47 10.91 37.05 29.62
Finch-3B 44.65 49.71 53.34 38.78 0.05 27.05 23.62 37.02 35.31 16.82 0.02 0.05 18.13 13.85 45.58 64.28 54.45 39.26 14.60 45.89 29.65
Finch-7B 52.17 49.41 48.60 32.05 3.45 38.68 27.42 42.19 35.50 19.04 0.66 10.19 27.33 21.50 60.03 60.44 64.57 45.26 23.19 60.19 34.39
Finch-14B 61.52 52.74 50.18 32.59 4.39 40.76 31.58 45.43 34.83 20.70 0.33 3.45 31.60 24.93 63.48 64.23 74.58 51.98 25.65 64.12 37.13
Gemma-2B 44.13 47.54 45.09 47.72 13.58 25.46 22.03 33.03 35.46 14.16 20.93 18.34 24.91 14.65 47.76 65.43 62.88 43.28 16.20 46.72 34.81
Gemma-7B 54.37 26.78 28.30 41.17 18.20 40.64 31.58 45.09 41.48 21.33 25.65 26.22 30.34 25.21 63.07 57.60 85.23 54.14 26.45 55.36 39.77
Gemma-2B (IT) 50.82 51.90 51.82 58.94 2.84 38.68 29.01 45.85 41.52 22.52 28.61 18.74 29.97 8.14 26.48 38.62 55.55 38.06 8.67 26.69 31.55
Gemma-7B (IT) 65.04 61.48 47.03 56.20 12.03 59.00 29.13 46.88 42.01 31.98 20.91 19.89 28.98 12.11 39.10 53.60 78.65 55.74 12.83 39.14 38.84
Gemma1.1-2B (IT) 49.87 49.48 52.16 56.77 4.68 47.98 29.01 45.46 41.19 25.77 24.31 17.69 30.93 9.09 28.74 38.38 63.14 34.65 9.31 29.17 32.99
Gemma1.1-7B (IT) 64.33 27.11 38.18 48.82 9.32 61.69 34.15 51.05 41.91 44.18 22.45 22.20 33.78 16.26 51.42 61.03 85.70 58.96 17.34 52.03 40.29
Gemma2-2B 49.57 50.14 49.53 42.74 14.64 22.15 24.11 36.23 37.32 10.72 15.66 16.43 33.18 17.87 53.16 69.09 73.33 39.57 18.31 50.28 35.83
Gemma2-9B 68.56 52.74 47.54 34.40 21.35 31.82 30.35 45.57 42.01 13.12 17.02 25.73 36.65 28.67 66.26 67.82 91.55 62.06 25.48 65.34 43.55
Gemma2-27B 75.94 32.38 42.14 35.68 24.81 40.76 33.29 47.81 42.49 17.56 25.24 32.61 36.52 36.84 78.42 66.29 93.38 68.28 38.56 77.05 47.44
Gemma2-2B (IT) 59.45 34.61 46.97 48.82 11.43 43.21 36.96 53.15 44.78 26.43 15.05 16.10 37.71 17.12 49.70 68.30 85.51 58.53 17.89 45.90 38.96
Gemma2-9B (IT) 73.18 60.80 69.62 54.06 16.80 41.74 42.84 60.11 46.99 53.60 16.64 19.41 40.05 25.18 63.93 64.76 91.41 68.53 23.27 66.04 49.00
Gemma2-27B (IT) 79.85 65.31 73.38 48.50 20.28 43.94 46.39 64.48 50.14 71.90 26.02 30.35 40.23 34.32 77.52 60.67 93.11 69.28 36.23 77.22 55.38
RecurrentGemma-2B 49.28 35.29 47.21 54.74 7.64 25.70 21.05 35.10 37.94 14.12 21.50 17.32 27.29 13.41 47.11 67.54 69.06 42.97 15.10 46.34 33.26
RecurrentGemma-9B 45.97 36.14 44.85 42.38 16.01 33.78 26.32 38.37 37.28 18.96 22.25 21.73 30.76 24.35 64.86 63.68 84.11 51.25 26.84 63.77 40.16
RecurrentGemma-2B (IT) 46.69 35.76 48.57 51.28 5.75 33.78 26.44 42.77 39.62 21.81 22.65 19.88 30.05 11.14 36.24 67.22 77.81 46.60 11.55 36.37 34.71
RecurrentGemma-9B (IT) 60.28 18.42 44.47 54.98 18.77 45.53 32.19 47.60 43.31 32.27 23.62 24.05 34.12 21.86 60.37 55.99 87.25 64.61 23.52 60.69 41.63
LLaMA2-7B 49.96 47.01 44.12 39.21 20.01 33.05 25.09 38.99 39.52 18.67 23.66 26.73 29.00 26.65 64.64 54.45 72.59 51.38 28.98 63.75 38.65
LLaMA2-13B 73.00 68.88 43.50 34.58 21.43 29.50 26.07 36.90 39.48 16.56 23.92 28.74 32.61 31.00 70.08 66.45 83.17 62.35 33.68 68.54 43.95
LLaMA2-70B 69.12 51.44 46.76 42.74 31.90 38.80 31.09 44.86 42.97 17.74 24.38 31.63 33.77 37.37 80.07 56.91 87.49 66.63 40.83 77.88 47.43
LLaMA2-7B (IT) 58.17 49.06 45.51 50.53 7.03 42.96 29.87 44.62 44.02 25.32 27.80 23.97 33.01 23.63 57.40 56.34 80.56 57.26 24.74 55.80 40.69
LLaMA2-13B (IT) 67.62 58.38 48.70 42.35 9.63 42.11 28.03 43.96 47.12 32.24 27.60 28.60 35.97 27.01 66.44 64.75 85.16 60.71 29.22 65.57 43.13
LLaMA2-70B (IT) 97.26 63.49 60.52 55.88 12.95 48.35 35.74 52.75 43.92 38.82 24.73 30.17 35.75 32.22 73.76 58.52 85.83 65.18 35.29 71.47 49.37
LLaMA3-8B 72.04 34.85 46.92 33.76 16.56 40.15 27.05 43.96 40.29 10.35 24.51 28.72 33.61 30.19 71.59 61.88 88.12 60.84 32.69 70.26 42.62
LLaMA3-70B 77.09 70.55 51.85 33.01 26.02 36.47 29.62 45.72 41.82 10.17 24.98 32.84 40.70 42.41 81.87 65.56 93.84 69.57 44.04 78.39 50.15
LLaMA3-8B (IT) 76.01 68.15 64.35 59.19 15.81 47.86 35.99 51.60 46.03 38.45 24.56 25.49 37.41 27.51 67.56 56.97 88.94 65.66 28.98 67.49 48.66
LLaMA3-70B (IT) 82.79 88.26 72.77 41.88 27.67 50.31 44.06 61.83 46.99 52.50 24.40 30.06 40.23 36.45 80.20 65.02 93.53 70.56 38.56 72.89 54.75
Jamba 56.70 30.10 50.05 69.34 0.45 55.94 28.76 45.39 41.53 17.38 22.93 24.52 27.89 25.57 58.29 67.82 83.63 59.88 33.16 68.47 42.73
Jamba (IT) 55.11 20.45 49.32 74.47 0.00 64.99 32.93 49.42 44.98 0.74 7.34 9.13 32.20 19.75 50.98 63.94 85.00 57.28 27.51 62.57 38.68
Mamba-130M 48.08 56.71 48.36 71.58 1.93 29.62 23.01 41.75 27.18 17.01 11.87 13.81 1.43 1.44 3.58 72.14 50.91 33.84 1.58 3.59 27.11
Mamba-370M 47.53 42.29 49.48 59.94 1.60 26.93 22.64 38.22 29.67 18.67 14.68 15.41 5.49 2.80 12.33 71.29 51.49 38.55 3.71 13.80 27.15
Mamba-790M 43.22 39.69 49.74 50.00 7.54 28.03 21.91 35.78 33.01 20.70 17.88 16.04 7.09 4.93 20.08 68.95 52.44 38.73 5.96 20.88 28.51
Mamba-1.4B 44.07 39.72 51.36 47.54 8.51 24.85 20.81 35.21 34.26 18.67 17.95 16.28 8.95 7.48 26.84 70.90 53.34 39.08 7.89 28.25 29.53
Mamba-2.8B 49.54 48.03 49.96 47.33 11.90 24.85 22.15 35.71 33.78 20.15 19.41 20.60 13.80 10.50 36.70 61.77 56.19 42.99 12.19 37.30 31.89
Mamba2-130M 51.32 46.95 49.60 71.30 1.62 33.78 23.26 43.17 29.61 20.96 12.49 15.27 3.03 1.50 4.03 72.52 49.75 39.17 1.55 3.53 27.09
Mamba2-370B 45.84 49.71 49.89 54.06 1.30 31.82 23.75 41.44 31.00 23.11 17.80 16.30 7.82 3.30 11.46 69.79 51.83 39.08 3.93 12.51 28.79
Mamba2-780B 47.88 44.28 51.65 52.39 3.44 26.56 23.38 38.02 33.73 20.78 20.93 17.36 12.70 4.35 18.87 69.78 51.27 38.35 4.74 19.98 28.92
Mamba2-1.3B 50.48 47.19 49.97 48.93 5.26 23.50 20.69 35.98 36.08 20.33 18.78 17.19 14.42 7.37 27.12 69.00 52.23 41.26 8.34 27.39 30.36
Mamba2-2.7B 50.49 47.23 46.04 40.06 10.56 21.05 20.69 33.50 35.89 22.74 21.59 19.73 17.26 11.02 36.27 68.69 55.66 40.35 12.99 36.47 32.33
Mamba2Attn-2.7B 41.48 34.69 50.71 41.77 10.16 26.19 20.20 32.95 37.70 17.38 17.91 10.02 11.62 6.01 21.76 79.51 54.64 38.87 6.18 21.76 32.13
Mistral-7B 59.50 52.69 50.12 39.96 19.64 40.76 27.91 42.53 40.77 16.82 24.87 28.67 33.77 29.11 69.92 54.42 85.65 58.59 31.33 68.71 42.73
Mistral-7B (IT) 69.87 41.12 62.24 45.94 18.77 56.67 42.11 59.65 46.70 44.55 23.59 28.40 35.59 28.45 68.35 56.97 89.01 65.37 29.81 67.77 47.64
Mixtral-8X7B 66.62 69.32 77.13 36.11 31.89 43.33 34.39 48.65 39.90 17.56 25.47 30.35 36.13 37.40 79.17 63.21 91.27 68.35 40.00 77.09 50.44
Mixtral-8X7B (IT) 81.17 73.47 62.52 42.31 28.53 57.04 49.82 64.84 46.70 46.58 25.31 28.38 35.69 35.54 77.55 58.55 91.09 68.84 37.06 76.18 53.10
Pythia-160M 51.59 45.57 51.35 72.44 0.02 52.51 24.72 46.48 26.99 8.69 3.29 2.72 6.06 0.11 0.28 50.32 50.16 37.96 0.11 0.28 20.00
Pythia-410M 51.72 48.36 48.85 64.10 1.61 28.89 22.77 41.03 31.20 16.45 15.52 14.88 9.11 1.97 9.11 69.94 50.12 39.49 1.97 9.11 25.46
Pythia-1B 45.27 51.59 49.82 54.38 2.47 27.66 23.99 40.48 32.63 17.38 18.95 15.58 11.42 2.77 14.07 72.05 50.34 39.22 2.77 14.07 26.50
Pythia-1.4B 46.86 48.50 52.36 56.62 4.68 25.95 22.89 39.10 34.45 16.27 19.42 16.53 13.18 5.18 18.70 71.74 51.98 39.22 6.40 20.33 29.96
Pythia-2.8B 45.00 42.44 55.38 50.96 6.00 23.62 21.05 36.27 34.74 16.45 20.99 19.39 20.05 7.95 27.41 70.95 51.98 44.43 9.06 30.11 30.86
Pythia-6.9B 47.58 48.24 44.98 47.23 5.57 25.70 21.66 36.32 37.83 20.01 24.13 22.15 24.76 8.70 34.43 70.39 51.88 42.91 11.41 35.97 30.52
Pythia-12B 49.65 45.29 54.94 42.41 7.73 25.70 20.56 32.04 37.99 18.11 22.72 20.77 21.73 10.61 39.77 67.31 54.41 42.14 12.69 41.60 32.98
Table 3: Complete results on individual tasks.
Hallu. Detection Instr. Following QA Reading Comp. Sum. Fact-Checking
Falcon-7B 52.40 25.54 33.03 29.18 20.70 46.91
Falcon-7B (IT) 49.50 38.16 28.61 25.54 19.70 39.20
FalconMamba-7B 55.80 42.97 39.94 23.76 23.89 61.85
FalconMamba-7B (IT) 56.53 41.78 39.51 22.91 23.80 60.05
Gemma-2B 53.01 30.94 27.43 30.35 19.63 43.28
Gemma-7B 50.45 31.25 38.20 39.96 25.93 54.14
Gemma-2B (IT) 49.74 40.73 23.30 27.52 23.67 38.06
Gemma-7B (IT) 61.16 44.09 31.28 25.00 20.40 55.74
Gemma1.1-2B (IT) 50.61 41.27 25.43 29.38 21.00 34.65
Gemma1.1-7B (IT) 55.27 46.50 36.66 26.63 22.32 58.96
Gemma2-2B 58.33 26.73 29.59 32.95 16.04 39.57
Gemma2-9B 65.64 23.76 39.36 39.74 21.38 62.06
Gemma2-27B 62.03 26.62 47.19 43.04 28.92 68.28
Gemma2-2B (IT) 58.97 37.63 34.42 27.69 15.57 58.53
Gemma2-9B (IT) 71.95 53.83 42.49 39.04 18.03 68.53
Gemma2-27B (IT) 74.46 60.20 50.05 48.06 28.18 69.28
RecurrentGemma-2B 53.68 34.43 27.36 31.66 19.41 42.97
RecurrentGemma-9B 54.95 30.67 36.79 35.57 21.99 51.25
RecurrentGemma-2B (IT) 55.21 36.55 25.51 28.89 21.27 46.60
RecurrentGemma-9B (IT) 53.28 43.63 38.82 28.91 23.83 64.61
LLaMA2-7B 53.63 28.94 37.64 27.58 25.19 51.38
LLaMA2-13B 67.00 25.57 39.65 31.17 26.33 62.35
LLaMA2-70B 62.34 30.24 47.85 39.48 28.00 66.63
LLaMA2-7B (IT) 57.93 37.93 35.75 31.32 25.88 57.26
LLaMA2-13B (IT) 64.92 37.29 39.00 29.53 28.10 60.71
LLaMA2-70B (IT) 73.12 47.35 45.32 31.27 27.45 65.18
LLaMA3-8B 60.76 22.06 41.55 33.52 26.62 60.84
LLaMA3-70B 71.78 21.59 48.07 46.38 28.91 69.57
LLaMA3-8B (IT) 70.88 48.82 42.85 37.14 25.02 65.66
LLaMA3-70B (IT) 80.47 47.19 51.50 38.65 27.23 70.56
Jamba 57.66 43.36 39.50 33.00 23.72 59.88
Jamba (IT) 54.76 37.60 38.52 27.12 8.24 57.28
Finch-1.6B 53.09 31.10 22.69 16.60 11.77 39.47
Finch-3B 53.29 27.80 25.96 17.89 0.03 39.26
Finch-7B 55.04 25.55 34.58 21.05 5.42 45.26
Finch-14B 60.65 26.64 37.54 22.40 1.89 51.98
Mamba-130M 55.24 44.29 13.31 12.87 12.84 33.84
Mamba-370M 52.41 39.30 15.25 12.97 15.04 38.55
Mamba-790M 50.81 35.35 18.14 18.75 16.96 38.73
Mamba-1.4B 51.88 33.11 19.98 20.44 17.11 39.08
Mamba-2.8B 53.10 33.74 23.91 20.80 20.00 42.99
Mamba2-130M 54.02 46.13 14.05 12.73 13.88 39.17
Mamba2-370B 53.41 38.58 16.19 19.26 17.05 39.08
Mamba2-780B 52.97 36.58 17.42 21.44 19.15 38.35
Mamba2-1.3B 53.77 34.63 19.45 22.20 17.98 41.26
Mamba2-2.7B 53.62 31.40 22.82 27.90 20.66 40.35
Mamba2Attn-2.7B 53.60 29.57 19.64 36.83 21.72 43.37
Mistral-7B 60.48 28.39 41.24 32.03 26.77 58.59
Mistral-7B (IT) 63.84 45.24 46.45 33.95 26.00 65.37
Mixtral-8X7B 73.51 26.84 48.99 40.66 27.91 68.35
Mixtral-8X7B (IT) 73.36 44.44 53.32 36.76 26.84 68.84
Pythia-160M 49.80 40.56 10.42 8.14 3.00 37.96
Pythia-410M 53.18 40.28 11.72 20.58 15.20 39.49
Pythia-1B 53.81 35.88 13.78 27.11 17.26 39.22
Pythia-1.4B 54.29 36.45 17.90 22.16 17.98 39.22
Pythia-2.8B 53.15 33.71 20.18 22.89 20.19 44.43
Pythia-6.9B 52.62 33.62 22.47 24.86 23.14 42.91
Pythia-12B 54.32 30.26 23.84 28.05 21.75 42.14
Table 4: Results separated by task categories.
Faithfulness Factuality
Falcon-7B 34.91 36.92
Falcon-7B (IT) 34.20 33.17
FalconMamba-7B 35.92 47.02
FalconMamba-7B (IT) 35.66 46.64
Finch-1.6B 31.62 27.41
Finch-3B 29.21 30.14
Finch-7B 30.52 38.65
Finch-14B 32.08 42.69
Gemma-2B 35.85 32.56
Gemma-7B 36.48 44.50
Gemma-2B (IT) 36.77 28.00
Gemma-7B (IT) 39.19 38.46
Gemma1.1-2B (IT) 36.60 30.12
Gemma1.1-7B (IT) 37.11 43.79
Gemma2-2B 36.61 34.96
Gemma2-9B 40.56 46.85
Gemma2-27B 41.54 53.92
Gemma2-2B (IT) 36.26 41.94
Gemma2-9B (IT) 48.11 49.98
Gemma2-27B (IT) 54.56 56.28
RecurrentGemma-2B 36.55 31.46
RecurrentGemma-2B (IT) 36.41 32.85
RecurrentGemma-9B 36.61 42.96
RecurrentGemma-9B (IT) 36.44 46.24
Jamba 39.80 45.95
Jamba (IT) 32.89 45.04
LLaMA2-7B 35.14 42.51
LLaMA2-13B 40.83 46.27
LLaMA2-70B 41.74 53.69
LLaMA2-7B (IT) 39.15 42.39
LLaMA2-13B (IT) 41.71 45.78
LLaMA2-70B (IT) 47.56 51.35
LLaMA3-8B 37.60 48.14
LLaMA3-70B 45.92 54.80
LLaMA3-8B (IT) 47.69 49.74
LLaMA3-70B (IT) 52.15 57.61
Mamba-130M 34.38 19.13
Mamba-370M 32.56 21.21
Mamba-790M 32.95 23.63
Mamba-1.4B 33.44 25.23
Mamba-2.8B 34.47 29.05
Mamba2-130M 34.42 20.13
Mamba2-370B 34.93 22.04
Mamba2-780B 35.40 22.90
Mamba2-1.3B 35.31 24.91
Mamba2-2.7B 36.39 27.86
Mamba2Attn-2.7B 38.91 26.49
Mistral-7B 38.47 47.42
Mistral-7B (IT) 43.14 52.59
Mixtral-8X7B 46.16 55.15
Mixtral-8X7B (IT) 48.05 58.65
Pythia-160M 25.13 15.23
Pythia-410M 33.29 18.69
Pythia-1B 35.43 18.21
Pythia-1.4B 35.89 23.44
Pythia-2.8B 35.48 25.79
Pythia-6.9B 36.30 27.46
Pythia-12B 36.85 28.73
Table 5: Results categorized by faithfulness and factuality.
B.2 Instruction Tuning Changes
Figure 5 and Figure 6 contain the numerical values for changes in task performance from base models to instruction-tuned models.
Figure 5: Change in task performance from base model to instruction fine-tuned model, for all tasks.
Figure 6: Change in task category performance from base model to instruction fine-tuned model.